A 1.00-mm-diameter glass sphere has a charge of + 1.30nC.
What speed does an electron need to orbit the sphere 1.50mm above the surface?

Answer should be in m/s. Thanks:)
Answer #1
Radius of electronsorbit r = heightof electron abovesurface + radius of sphere
= 1.50mm + 1.00 /2 mm
= 2.00 mm
= 2.00* 10-3 m
Force exerted on theelectron F = (1/4πε0)* q * e / r2
q = chargeon sphere
F = 9.0 *109 * 1.30 * 10-9 * 1.6 * 10-19 /(2.00 * 10-3)2
= 4.68* 10-13 N
--------------------------
This force works as centripetal force,hence
F = m* v2 / r
v2 = F* r / m
    = 4.68* 10-13 * 2.00 * 10-3 / 9.1 *10-31
    = 1.028* 1015
required velocity ofelectron v = √(10.28* 1014)
=3.20* 107 m/s